Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Jerome's refrigerator contains 4 cans of cola

Expert replies
by BTGmoderatorDC » Thu Jan 04, 2018 6:44 am
Jerome's refrigerator contains 4 cans of cola, 4 cans of root beer, and 4 cans of ginger ale. If Jerome were to randomly pull two cans out of his refrigerator, what is the probability that he would pull out one can of cola and one can of root beer?

A. 1/11

B. 1/9

C. 8/33

D. 3/11

E. 1/3

Can some experts show the best solution?

OA C
Join the discussion
Source: — Problem Solving |

by DavidG@VeritasPrep » Thu Jan 04, 2018 7:15 am
lheiannie07 wrote:Jerome's refrigerator contains 4 cans of cola, 4 cans of root beer, and 4 cans of ginger ale. If Jerome were to randomly pull two cans out of his refrigerator, what is the probability that he would pull out one can of cola and one can of root beer?

A. 1/11

B. 1/9

C. 8/33

D. 3/11

E. 1/3

Can some experts show the best solution?

OA C
P(x) = # ways x can occur/# Total possibilities;

# Total possibilities: There are 12 cans, and we're selecting 2, so we want 12C2 = 12*11/2! = 66. (We can select any of the 12 cans with our first selection and any of the 11 remaining cans with our second selection. We divide by 2! because the order doesn't matter.)

# ways we can select a can of cola = 4. # ways we can select a can of root beer = 4. # ways we can select a can of cola and one can of root beer = 4*4 = 16.

16/66 = 8/33. The answer is C
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by DavidG@VeritasPrep » Thu Jan 04, 2018 7:19 am
lheiannie07 wrote:Jerome's refrigerator contains 4 cans of cola, 4 cans of root beer, and 4 cans of ginger ale. If Jerome were to randomly pull two cans out of his refrigerator, what is the probability that he would pull out one can of cola and one can of root beer?

A. 1/11

B. 1/9

C. 8/33

D. 3/11

E. 1/3

Can some experts show the best solution?

OA C
Alternatively, we can go selection by selection. Say we pick a root beer can followed by a cola can

P(selecting root beer first) = 4/12 = 1/3
P(selecting cola after having selected a root beer) = 4/11
P(selecting root beer first) * P(selecting cola after having selected a root beer) = 1/3 * 4/11 = 4/33

Of course, we could also have selected the cola first. The math will end up being identical:
P(selecting cola first) = 4/12 = 1/3
P(selecting root beer after having selected a cola) = 4/11
P(selecting cola first) *P(selecting root beer after having selected a cola) = 1/3 * 4/11 = 4/33

Now add 'em up: 4/33 + 4/33 = 8/33. The answer is C
Veritas Prep | GMAT Instructor

Veritas Prep Reviews
Save $100 off any live Veritas Prep GMAT Course
Join the discussion

by Brent@GMATPrepNow » Thu Jan 04, 2018 8:24 am
lheiannie07 wrote:Jerome's refrigerator contains 4 cans of cola, 4 cans of root beer, and 4 cans of ginger ale. If Jerome were to randomly pull two cans out of his refrigerator, what is the probability that he would pull out one can of cola and one can of root beer?

A. 1/11
B. 1/9
C. 8/33
D. 3/11
E. 1/3
P(one can of cola and one can of root beer) = P(1st can is cola AND 2nd can is root beer OR 1st can is root beer AND 2nd can is cola)
= P(1st can is cola AND 2nd can is root beer) + P(1st can is root beer AND 2nd can is cola)
= P(1st can is cola) x 2nd can is root beer) + P(1st can is root beer) x P(2nd can is cola)
= 4/12 x 4/11 + 4/12 x 4/11
= 4/33 + 4/33
= 8/33
= C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion